OpenStudy (anonymous):

4sin(theta)+3sqrt3=sqrt3 solve w/ interval 0 < theta < 2pi

OpenStudy (anonymous):

Pretty straightforward \[4 \sin (\theta) = \sqrt{3} - 3\sqrt{3}\]

OpenStudy (anonymous):

\[4\sin (\theta)=-2\sqrt{3}\]

OpenStudy (anonymous):

\[\sin (\theta)=-2\sqrt{3 } / 4\]

OpenStudy (anonymous):

\[\sin (\theta)=-\sqrt{3} / 2\]

OpenStudy (anonymous):

\[\theta = \sin^{-1} (-\sqrt{3}/2)\]

OpenStudy (anonymous):

sin THETA is NEGATIVE only in 3rd and 4th quadrant

OpenStudy (anonymous):

\[\theta=\sin^{-1} (\sin(\pi+\pi/3))\]

OpenStudy (anonymous):

THETA = 4 PI /3

OpenStudy (anonymous):

OR THETA = 2PI - PI/3 = 5PI/3

OpenStudy (anonymous):

TWO ANSWERS as the interval is from 0 -> 2 Pi